A flexible rectangular coil of wire with 150 loops is stretched in such a way that its dimension changes from 5 cm x 8 cm to 7 cm x 11 cm in 0.15 s in a magnetic field of 25T that is 30 degrees relative to the plane of the coil.

a. Calculate the induced EMF

b. How much energy in joules was dissipated in the circuit if the total resistance is 100Ω?

Show full solution:


1
Expert's answer
2022-05-20T04:53:16-0400

"\varepsilon=\frac{n∆SB\cos\alpha}t=46.3~V,"

"W=\frac{\varepsilon^2}R=21.4~J."